# Joint Beamforming and Antenna Position Optimization for Fluid Antenna-Assisted MU-MIMO Networks

### v1

@@ III-D The Update Step of $\mathbf{T}$
Unlike the update steps for $\underline{\mathbf{\Gamma}}$, $\underline{\mathbf{\Phi}}$, and $\underline{\mathbf{W}}$, the objective function $f_{\rm Quad}\left(\mathbf{T}\right)$ remains nonconcave w.r.t. $\mathbf{T}$. The MM algorithm [36] effectively solves this non-concave problem by iteratively finding a series of concave lower bounds for the non-concave function $f_{\rm Quad}\left(\mathbf{T}\right)$, known as the surrogate function. A key advantage of the MM algorithm is its convergence guarantee, which is discussed in the sequel.
To solve the problem (21) using the MM algorithm, the optimal value of $\mathbf{T}$ is computed iteratively. Each MM iteration consists of a majorization step, followed by a maximization step. In the majorization step, the surrogate function is computed as
@@
where $\mathtt{Tx}$ indicates the update step is related to $\mathbf{T}$, and $\overline{\mathbf{T}}$ denotes a constant satisfying (12c) and (12e). In the maximization step, we determine the optimal $\mathbf{T}$ subject to the given constraints as
@@
In each MM iteration, after constructing a distinct surrogate function, we obtain a new optimum accordingly. Then, the update step of $\mathbf{T}$ for the proposed MM algorithm is executed until convergence. The convergence of MM method is ensured because $f_{\rm Quad}(\mathbf{T})$ remains non-decreasing after each MM iteration, as
where $\mathbf{T}$ is the solution to the problem (23). The first inequality follows from (22), the second inequality holds since $\mathbf{T}$ is the optimal solution to the problem (23), and the last equality results from the equality condition of (22).
The key to applying the MM algorithm is constructing a suitable surrogate function. Here, we introduce the following lemma to construct the surrogate function that locally approximates the objective function [36].
@@
By directly applying Lemma 1 to $f_{\rm Quad}\left(\mathbf{T}\right)$ and letting $\mathbf{L}=\delta^{\mathtt{Tx}}\mathbf{I}$ with
@@
for any given $\mathbf{T}$ satisfying (12c) and (12e), we can construct the surrogate function $h^{\mathtt{Tx}}\left(\mathbf{T}\big{|}\overline{\mathbf{T}}\right)$ as
@@
$h^{\mathtt{Tx}}\left(\mathbf{T}\big{|}\overline{\mathbf{T}}\right)=-\frac{\delta^{\mathtt{Tx}}}{2}\mathrm{vec}\left(\mathbf{T}\right)^{\mathsf{T}}\mathrm{vec}\left(\mathbf{T}\right)\\ +\left(\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}^{\mathsf{T}}\left(\overline{\mathbf{T}}\right)+\delta^{\mathtt{Tx}}\mathrm{vec}\left(\overline{\mathbf{T}}\right)^{\mathsf{T}}\right)\mathrm{vec}\left(\mathbf{T}\right)+\mathrm{const},$ (27)
where the entries of $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ are given by
$\frac{\partial f_{\rm Quad}}{\partial x_{m}^{t}}$ $=-\frac{4\pi}{\lambda}\sum_{k=1}^{K}\sum_{q=1}^{L_{k}^{t}}\Big{\lvert}[\mathbf{D}_{k}^{\mathtt{Tx}}]_{mq}\Big{\rvert}\cos\theta_{kq}^{t}\cos\phi_{kq}^{t}\sin\left(\xi_{kmq}^{\mathtt{Tx}}\right),$ (28a)
$\frac{\partial f_{\rm Quad}}{\partial y_{m}^{t}}$ $=-\frac{4\pi}{\lambda}\sum_{k=1}^{K}\sum_{q=1}^{L_{k}^{t}}\Big{\lvert}[\mathbf{D}_{k}^{\mathtt{Tx}}]_{mq}\Big{\rvert}\cos\theta_{kq}^{t}\sin\phi_{kq}^{t}\sin\left(\xi_{kmq}^{\mathtt{Tx}}\right),$ (28b)
$\frac{\partial f_{\rm Quad}}{\partial z_{m}^{t}}$ $=-\frac{4\pi}{\lambda}\sum_{k=1}^{K}\sum_{q=1}^{L_{k}^{t}}\Big{\lvert}[\mathbf{D}_{k}^{\mathtt{Tx}}]_{mq}\Big{\rvert}\sin\theta_{kq}^{t}\sin\left(\xi_{kmq}^{\mathtt{Tx}}\right),$ (28c)
respectively. Here, $\mathbf{D}_{k}^{\mathtt{Tx}}$ and $\xi_{kmq}^{\mathtt{Tx}}$ are given by
@@
$\mathbf{D}_{k}^{\mathtt{Tx}}\triangleq\left(\frac{\partial f_{\rm Quad}}{\partial\mathbf{G}_{k}}\right)^{\mathsf{T}}=\sqrt{\alpha_{k}}\overline{\mathbf{W}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\\ -\hat{\mathbf{W}}\mathbf{G}_{k}^{\mathsf{H}}\left(\overline{\mathbf{T}}\right)\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}$ (29)
and
@@
$\xi_{kmq}^{\mathtt{Tx}}=$ $\angle[\mathbf{D}_{k}^{\mathtt{Tx}}]_{mq}+\frac{2\pi}{\lambda}\left(x_{m}^{t}\cos\theta_{kq}^{t}\cos\phi_{kq}^{t}\right.$
$\left.+y_{m}^{t}\cos\theta_{kq}^{t}\sin\phi_{kq}^{t}+z_{m}^{t}\sin\theta_{kq}^{t}\right),$ (30)
respectively, where we define $\overline{\mathbf{F}}_{k}\triangleq\mathbf{F}_{k}\left(\overline{\mathbf{R}}_{k}\right)$ and $\hat{\mathbf{W}}\triangleq\sum_{k=1}^{K}\overline{\mathbf{W}}_{k}\overline{\mathbf{W}}_{k}^{\mathsf{H}}$. Moreover, by deriving an upper bound for the right hand side (r.h.s.) of (26) for any $\mathbf{T}$ satisfying (12c) and (12e), we obtain the closed-form expression of $\delta^{\mathtt{Tx}}$ as given by (31) at the bottom of the page. The detailed derivations of (28) and (31) are shown in Appendix -A.
Although we have found the concave surrogate function (27), the constraint (12e) still makes the optimization problem (23) nonconcave. Therefore, we apply Cauchy-Schwarz inequality to the left hand side (l.h.s.) of the constraint (12e) and obtain a linear constraint
@@
$\frac{\left(\overline{\mathbf{t}}_{m}-\overline{\mathbf{t}}_{m^{\prime}}\right)\left(\mathbf{t}_{m}-\mathbf{t}_{m^{\prime}}\right)^{\top}}{\lVert\overline{\mathbf{t}}_{m}-\overline{\mathbf{t}}_{m^{\prime}}\rVert_{2}}\geq D,\ \forall m\neq m^{\prime}.$ (32)
Thus, the original problem (21) can be formulated as
@@
which is a concave quadratic programming problem. As demonstrated in [23], solving (33) becomes simpler by initially assuming all constraints are inactive. This transformation simplifies the problem to an unconstrained quadratic optimization, whose closed-form solution is given by
@@
III-E The Update Step of $\underline{\mathbf{R}}$
$\delta^{\mathtt{Rx}}=\underset{1\leq n\leq N}{\max}\frac{24\pi^{2}}{\lambda^{2}}L_{k}^{r}\left(\left(\sum_{j=1}^{N}\Big{\lvert}[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)[\overline{\mathbf{\Phi}}_{k}]_{j}^{\mathsf{H}}\Big{\rvert}+\sqrt{N}\Big{\lVert}[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\Big{\rVert}_{2}\right)\lVert\mathbf{\Sigma}_{k}\overline{\mathbf{G}}_{k}\hat{\mathbf{W}}\overline{\mathbf{G}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\rVert_{2}+\sqrt{\frac{\alpha_{k}}{L_{k}^{r}}}\Big{\lVert}[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{W}}_{k}^{\mathsf{H}}\overline{\mathbf{G}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big{\rVert}_{2}\right).$ (43)
@@ III-E The Update Step of $\underline{\mathbf{R}}$
$f_{\rm Quad}\left(\underline{\mathbf{R}}\right)=\sum_{k=1}^{K}f_{\rm Quad}\left(\mathbf{R}_{k}\right).$ (35)
Since the terms of the r.h.s. of (35) do not couple with each other, it is feasible to optimize $f_{\rm Quad}\left(\mathbf{R}_{k}\right)$ independently. Therefore, we simply provide the update step of $\mathbf{R}_{k}$ in the remainder of this subsection. Then, the optimization problem is reformulated as
@@
$\underset{\mathbf{\underline{R}}}{\max}$ $h^{\mathtt{Rx}}_{k}\left(\mathbf{R}_{k}\big{|}\overline{\mathbf{R}}_{k}\right)$ (37a)
@@
$\frac{\left(\overline{\mathbf{r}}_{kn}-\overline{\mathbf{r}}_{kn^{\prime}}\right)\left(\mathbf{r}_{kn}-\mathbf{r}_{kn^{\prime}}\right)^{\top}}{\lVert\overline{\mathbf{r}}_{kn}-\overline{\mathbf{r}}_{kn^{\prime}}\rVert_{2}}\geq D,\ \forall n\neq n^{\prime},$ (37b)
where $\mathtt{Rx}$ indicates the update step relates to $\underline{\mathbf{R}}$. The function $h^{\mathtt{Rx}}_{k}\left(\mathbf{R}_{k}\big{|}\overline{\mathbf{R}}_{k}\right)$ is the surrogate function of $f_{\rm Quad}\left(\mathbf{R}_{k}\right)$, given by
$h^{\mathtt{Rx}}_{k}\left(\mathbf{R}_{k}\big{|}\overline{\mathbf{R}}_{k}\right)=-\frac{\delta^{\mathtt{Rx}}_{k}}{2}\mathrm{vec}\left(\mathbf{R}_{k}\right)^{\mathsf{T}}\mathrm{vec}\left(\mathbf{R}_{k}\right)\\ +\left(\nabla_{\mathrm{vec}\left(\mathbf{R}_{k}\right)}f_{\rm Quad}^{\mathsf{T}}\left(\overline{\mathbf{R}}_{k}\right)+\delta^{\mathtt{Rx}}_{k}\mathrm{vec}\left(\overline{\mathbf{R}}_{k}\right)^{\mathsf{T}}\right)\mathrm{vec}\left(\mathbf{R}_{k}\right)\\ +\mathrm{const},$ (38)
@@
$\frac{\partial f_{\rm Quad}}{\partial x_{kn}^{r}}$ $=-\frac{4\pi}{\lambda}\sum_{k=1}^{K}\sum_{q=1}^{L_{k}^{r}}\Big{\lvert}[\mathbf{D}_{k}^{\mathtt{Rx}}]_{nq}\Big{\rvert}\cos\theta_{kq}^{r}\cos\phi_{kq}^{r}\sin\left(\xi_{knq}^{\mathtt{Rx}}\right),$ (40a)
$\frac{\partial f_{\rm Quad}}{\partial y_{kn}^{r}}$ $=-\frac{4\pi}{\lambda}\sum_{k=1}^{K}\sum_{q=1}^{L_{k}^{r}}\Big{\lvert}[\mathbf{D}_{k}^{\mathtt{Rx}}]_{nq}\Big{\rvert}\cos\theta_{kq}^{r}\sin\phi_{kq}^{r}\sin\left(\xi_{knq}^{\mathtt{Rx}}\right),$ (40b)
$\frac{\partial f_{\rm Quad}}{\partial z_{kn}^{r}}$ $=-\frac{4\pi}{\lambda}\sum_{k=1}^{K}\sum_{q=1}^{L_{k}^{r}}\Big{\lvert}[\mathbf{D}_{k}^{\mathtt{Rx}}]_{nq}\Big{\rvert}\sin\theta_{kq}^{r}\sin\left(\xi_{knq}^{\mathtt{Rx}}\right),$ (40c)
respectively. The expressions of $\mathbf{D}_{k}^{\mathtt{Rx}}$ and $\xi_{knq}^{\mathtt{Rx}}$ are given by
$\mathbf{D}^{\mathtt{Rx}}_{k}=$ $\sqrt{\alpha_{k}}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{W}}_{k}^{\mathsf{H}}\overline{\mathbf{G}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}$
$-\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\mathbf{F}_{k}^{\mathsf{H}}\left(\overline{\mathbf{R}}_{k}\right)\mathbf{\Sigma}_{k}\overline{\mathbf{G}}_{k}\hat{\mathbf{W}}\overline{\mathbf{G}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}.$ (41)
@@
$\xi_{knq}^{\mathtt{Rx}}=$ $\angle[\mathbf{D}^{\mathtt{Rx}}_{k}]_{nq}+\frac{2\pi}{\lambda}\left(x_{kn}^{r}\cos\theta_{kq}^{r}\cos\phi_{kq}^{r}\right.$
$\left.+y_{kn}^{r}\cos\theta_{kq}^{r}\sin\phi_{kq}^{r}+z_{kn}^{r}\sin\theta_{kq}^{r}\right),$ (42)
respectively, where we define $\overline{\mathbf{G}}_{k}\triangleq\mathbf{G}_{k}\left(\overline{\mathbf{T}}\right)$. The closed-form expression of $\delta^{\mathtt{Rx}}_{k}$ satisfying (39) is given by (43) at the bottom of the page. The derivations of (40) and (43) follow the same steps as (28) and (31), respectively. Hence, we omit them for brevity.

### v2

@@ IV-D Update Step of $\mathbf{T}$
Unlike the update steps for $\underline{\mathbf{\Gamma}}$, $\underline{\mathbf{\Phi}}$, and $\underline{\mathbf{W}}$, the objective function $f_{\rm Quad}\left(\mathbf{T}\right)$ remains non-concave w.r.t. $\mathbf{T}$. The MM algorithm [46] effectively solves this non-convex problem by iteratively finding a series of concave lower bounds for the non-concave function $f_{\rm Quad}\left(\mathbf{T}\right)$, known as the surrogate function. A key advantage of the MM algorithm is its convergence guarantee, which is discussed in Section IV-G.
To solve the problem (20) using the MM algorithm, the optimal value of $\mathbf{T}$ is computed iteratively. Each MM iteration consists of a majorization step, followed by a maximization step. In the majorization step, we construct the surrogate function such that
@@
where $\mathtt{Tx}$ indicates the update step is related to $\mathbf{T}$. The equality (21) holds when $\mathbf{T}=\overline{\mathbf{T}}$, i.e.,
$h^{\mathtt{Tx}}\left(\overline{\mathbf{T}}\big|\overline{\mathbf{T}}\right)=f_{\rm Quad}\left(\overline{\mathbf{T}}\right).$ (22)
In the maximization step, we determine the optimal $\mathbf{T}$ subject to the given constraints by solving the following convex optimization problem:
@@
In each MM iteration, a distinct surrogate function is constructed, based on which a new solution $\mathbf{T}$ is obtained by maximizing the resulting concave function. Then, the update step of $\mathbf{T}$ for the proposed MM algorithm is executed until convergence. The key to applying the MM algorithm is constructing a suitable surrogate function. Here, we introduce the following lemma to construct the surrogate function that locally approximates the objective function [46].
@@
We apply Lemma 1 to $f_{\rm Quad}\left(\mathbf{T}\right)$ and let $\mathbf{L}=\delta^{\mathtt{Tx}}\mathbf{I}$ with
@@
for any given $\mathbf{T}$ satisfying (11c) and (11e), where $\lambda_{\max}\left(\cdot\right)$ denotes the maximum eigenvalue of a matrix. Then, we can construct the surrogate function $h^{\mathtt{Tx}}\left(\mathbf{T}\big|\overline{\mathbf{T}}\right)$ as
@@
$h^{\mathtt{Tx}}\left(\mathbf{T}\big|\overline{\mathbf{T}}\right)=-\frac{\delta^{\mathtt{Tx}}}{2}\mathrm{vec}\left(\mathbf{T}\right)^{\mathsf{T}}\mathrm{vec}\left(\mathbf{T}\right)\\ +\left(\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}^{\mathsf{T}}\left(\overline{\mathbf{T}}\right)+\delta^{\mathtt{Tx}}\mathrm{vec}\left(\overline{\mathbf{T}}\right)^{\mathsf{T}}\right)\mathrm{vec}\left(\mathbf{T}\right)+\mathrm{const}.$ (26)
Finding the surrogate function $h^{\mathtt{Tx}}\left(\mathbf{T}\big|\overline{\mathbf{T}}\right)$ is thus equivalent to computing the gradient $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ and determining the constant $\delta^{\mathtt{Tx}}$. Since $\mathbf{T}$ influences $f_{\rm Quad}\left(\mathbf{T}\right)$ only through the matrix $\mathbf{G}_{k}$ as defined in (15), we compute the gradient $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ using the matrix chain rule. Let $\mathbf{D}_{k}^{\mathtt{Tx}}$ denote the transpose of the first-order derivative of $f_{\rm Quad}\left(\mathbf{T}\right)$ w.r.t. $\mathbf{G}_{k}$, and it is given by [47]
@@
$\mathbf{D}_{k}^{\mathtt{Tx}}$ $\triangleq\left(\frac{\partial f_{\rm Quad}}{\partial\mathbf{G}_{k}}\right)^{\mathsf{T}}$
$=\sqrt{\alpha_{k}}\overline{\mathbf{W}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}-\hat{\mathbf{W}}\mathbf{G}_{k}^{\mathsf{H}}\left(\overline{\mathbf{T}}\right)\hat{\mathbf{\Sigma}}_{k},$ (27)
where we define $\overline{\mathbf{F}}_{k}\triangleq\mathbf{F}_{k}\left(\overline{\mathbf{R}}_{k}\right)$, $\hat{\mathbf{W}}\triangleq\sum_{k=1}^{K}\overline{\mathbf{W}}_{k}\overline{\mathbf{W}}_{k}^{\mathsf{H}}$, and $\hat{\mathbf{\Sigma}}_{k}^{\mathtt{Tx}}\triangleq\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}$. Based on $\mathbf{D}_{k}^{\mathtt{Tx}}$, the entries of $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ can be computed as
$\frac{\partial f_{\rm Quad}}{\partial\mathbf{t}_{m}}$ $=\left[\frac{\partial f_{\rm Quad}}{\partial x_{m}},\frac{\partial f_{\rm Quad}}{\partial y_{m}},\frac{\partial f_{\rm Quad}}{\partial z_{m}}\right]^{\mathsf{T}}$
$=-\frac{4\pi}{\lambda}\sum_{k=1}^{K}\sum_{q=1}^{L_{k}^{\mathtt{Tx}}}\Big\lvert[\mathbf{D}_{k}^{\mathtt{Tx}}]_{mq}\Big\rvert\sin(\xi_{kmq}^{\mathtt{Tx}})\mathbf{g}_{kq}^{\mathtt{Tx}},$ (28)
where $\xi_{kmq}^{\mathtt{Tx}}$ is calculated by
$\xi_{kmq}^{\mathtt{Tx}}=$ $\angle[\mathbf{D}_{k}^{\mathtt{Tx}}]_{mq}+\frac{2\pi}{\lambda}\rho_{kq}^{\mathtt{Tx}}(\mathbf{t}_{m}).$ (29)
The detailed derivation of (IV-D) is shown in Appendix -A. According to (25), the constant $\delta^{\mathtt{Tx}}$ can be chosen by finding the upper bound of the maximum eigenvalue of the Hessian matrix $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}^{2}f_{\rm Quad}\left(\mathbf{T}\right)$. Since calculating the eigenvalues of the Hessian matrix is computationally expensive, we first leverage the matrix infinity norm to upper bound the maximum eigenvalue. Then, we calculate the upper bound of the matrix infinity norm for all possible $\mathbf{T}$. The detailed derivations of $\delta^{\mathtt{Tx}}$ is provided in Appendix -B and the result is given by
@@
$\delta^{\mathtt{Tx}}=$ $\underset{1\leq m\leq M}{\max}\frac{24\pi^{2}}{\lambda^{2}}\sum_{k=1}^{K}\left[\Big(\sum_{j=1}^{M}\Big\lvert[\hat{\mathbf{W}}]_{mj}\Big\rvert+\sqrt{M}\Big\lVert[\hat{\mathbf{W}}]_{m}\Big\rVert_{2}\Big)\lVert\hat{\mathbf{\Sigma}}_{k}^{\mathtt{Tx}}\rVert_{2}\right$.
@@
$+\left.\sqrt{\frac{\alpha_{k}}{L_{k}^{\mathtt{Tx}}}}\Big\lVert[\overline{\mathbf{W}}_{k}]_{m}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big\rVert_{2}\right]L_{k}^{\mathtt{Tx}}.$ (30)
Although we have found the concave surrogate function (26), the non-convex constraint (11e) still makes the optimization problem (23) non-convex. To make the problem tractable, we apply the Cauchy-Schwarz inequality to the left-hand side (l.h.s.) of the constraint (11e) and obtain a lower bound of $\lVert\mathbf{t}_{m}-\mathbf{t}_{m^{\prime}}\rVert_{2}$, given by
$\lVert\mathbf{t}_{m}-\mathbf{t}_{m^{\prime}}\rVert_{2}\geq\frac{\left(\overline{\mathbf{t}}_{m}-\overline{\mathbf{t}}_{m^{\prime}}\right)^{\mathsf{T}}\left(\mathbf{t}_{m}-\mathbf{t}_{m^{\prime}}\right)}{\lVert\overline{\mathbf{t}}_{m}-\overline{\mathbf{t}}_{m^{\prime}}\rVert_{2}},\ \forall m\neq m^{\prime}.$ (31)
The inequality (31) indicates that if $\mathbf{T}$ satisfies the following constraint
$\frac{\left(\overline{\mathbf{t}}_{m}-\overline{\mathbf{t}}_{m^{\prime}}\right)^{\mathsf{T}}\left(\mathbf{t}_{m}-\mathbf{t}_{m^{\prime}}\right)}{\lVert\overline{\mathbf{t}}_{m}-\overline{\mathbf{t}}_{m^{\prime}}\rVert_{2}}\geq D,\ \forall m\neq m^{\prime},$ (32)
then $\mathbf{T}$ also satisfies constraint (11e). In other words, constraint (32) is a sufficient condition for constraint (11e). Therefore, the problem (20) is iteratively solved by
@@
which is a convex quadratic programming problem. As demonstrated in [16], solving (33) can be simplified by initially assuming all constraints are inactive. This transformation reduces the problem to an unconstrained quadratic optimization, whose closed-form solution is given by
@@
Remark: Different from [15, 31], the surrogate function $h^{\mathtt{Tx}}(\mathbf{T}|\overline{\mathbf{T}})$ in this paper is constructed w.r.t. all transmit FAs $\mathbf{T}$, enabling the parallel update of all FA positions. Moreover, prior works [15, 31] employ the inequality from [46, Eq. (26)] to eliminate the second-order term w.r.t. FA positions. In contrast, we construct the surrogate function directly from the gradient and Hessian of $f_{\rm Quad}$, yielding a tighter approximation and improved optimization performance. This construction relies on matrix chain rules (-A) and the matrix infinity norm bound in (-B) detailed in Appendix -A and Appendix -B, respectively.
IV-E Update Step of $\underline{\mathbf{R}}$
$\delta^{\mathtt{Rx}}=\underset{1\leq n\leq N}{\max}\frac{24\pi^{2}}{\lambda^{2}}L_{k}^{\mathtt{Rx}}\left[\Big(\sum_{j=1}^{N}\Big\lvert[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)[\overline{\mathbf{\Phi}}_{k}]_{j}^{\mathsf{H}}\Big\rvert+\sqrt{N}\Big\lVert[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\Big\rVert_{2}\Big)\lVert\hat{\mathbf{\Sigma}}_{k}^{\mathtt{Rx}}\rVert_{2}+\sqrt{\frac{\alpha_{k}}{L_{k}^{\mathtt{Rx}}}}\Big\lVert[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{W}}_{k}^{\mathsf{H}}\overline{\mathbf{G}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big\rVert_{2}\right].$ (43)
@@ IV-E Update Step of $\underline{\mathbf{R}}$
$f_{\rm Quad}\left(\underline{\mathbf{R}}\right)=\sum\nolimits_{k=1}^{K}f_{\rm Quad}\left(\mathbf{R}_{k}\right).$ (35)
Since the terms of the right hand side (r.h.s.) of (35) do not couple with each other, it is feasible to optimize $f_{\rm Quad}\left(\mathbf{R}_{k}\right)$ independently. Therefore, we simply provide the update step of $\mathbf{R}_{k}$ in the remainder of this subsection. The optimization problem is then reformulated as
@@
$\underset{\mathbf{\underline{R}}}{\max}\ \$$h^{\mathtt{Rx}}_{k}\left(\mathbf{R}_{k}\big|\overline{\mathbf{R}}_{k}\right)$ (37a)
@@
$\frac{\left(\overline{\mathbf{r}}_{kn}-\overline{\mathbf{r}}_{kn^{\prime}}\right)^{\mathsf{T}}\left(\mathbf{r}_{kn}-\mathbf{r}_{kn^{\prime}}\right)}{\lVert\overline{\mathbf{r}}_{kn}-\overline{\mathbf{r}}_{kn^{\prime}}\rVert_{2}}\geq D,\ \forall n\neq n^{\prime},$ (37b)
where $\mathtt{Rx}$ indicates the update step relates to $\underline{\mathbf{R}}$. The function $h^{\mathtt{Rx}}_{k}\left(\mathbf{R}_{k}\big|\overline{\mathbf{R}}_{k}\right)$ is the surrogate function of $f_{\rm Quad}\left(\mathbf{R}_{k}\right)$:
$h^{\mathtt{Rx}}_{k}\left(\mathbf{R}_{k}\big|\overline{\mathbf{R}}_{k}\right)=-\frac{\delta^{\mathtt{Rx}}_{k}}{2}\mathrm{vec}\left(\mathbf{R}_{k}\right)^{\mathsf{T}}\mathrm{vec}\left(\mathbf{R}_{k}\right)\\ +\left(\nabla_{\mathrm{vec}\left(\mathbf{R}_{k}\right)}f_{\rm Quad}^{\mathsf{T}}\left(\overline{\mathbf{R}}_{k}\right)+\delta^{\mathtt{Rx}}_{k}\mathrm{vec}\left(\overline{\mathbf{R}}_{k}\right)^{\mathsf{T}}\right)\mathrm{vec}\left(\mathbf{R}_{k}\right)+\mathrm{const},$ (38)
@@
$\frac{\partial f_{\rm Quad}}{\partial\mathbf{r}_{kn}}$ $=-\frac{4\pi}{\lambda}\sum_{k=1}^{K}\sum_{q=1}^{L_{k}^{\mathtt{Rx}}}\Big\lvert[\mathbf{D}_{k}^{\mathtt{Rx}}]_{nq}\Big\rvert\sin(\xi_{knq}^{\mathtt{Rx}})\mathbf{f}_{kq}^{\mathtt{Rx}}.$ (40a)
The expressions of $\mathbf{D}_{k}^{\mathtt{Rx}}$ and $\xi_{knq}^{\mathtt{Rx}}$ are given by
$\mathbf{D}^{\mathtt{Rx}}_{k}=\left(\frac{\partial f_{\rm Quad}}{\partial\mathbf{F}_{k}}\right)^{\mathsf{T}}=$ $\sqrt{\alpha_{k}}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{W}}_{k}^{\mathsf{H}}\overline{\mathbf{G}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}$
$-\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\mathbf{F}_{k}^{\mathsf{H}}\left(\overline{\mathbf{R}}_{k}\right)\hat{\mathbf{\Sigma}}_{k}^{\mathtt{Rx}}$ (41)
@@
$\xi_{knq}^{\mathtt{Rx}}=\angle[\mathbf{D}^{\mathtt{Rx}}_{k}]_{nq}+\frac{2\pi}{\lambda}\rho_{kq}^{\mathtt{Rx}}(\mathbf{r}_{kn}),$ (42)
respectively, where we define $\overline{\mathbf{G}}_{k}\triangleq\mathbf{G}_{k}\left(\overline{\mathbf{T}}\right)$ and $\hat{\mathbf{\Sigma}}_{k}^{\mathtt{Rx}}=\mathbf{\Sigma}_{k}\overline{\mathbf{G}}_{k}\hat{\mathbf{W}}\overline{\mathbf{G}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}$. The closed-form expression of $\delta^{\mathtt{Rx}}_{k}$ satisfying (39) is given by (43) at the bottom of the page. The derivations of (40) and (43) follow the same steps as (IV-D) and (IV-D), respectively. Hence, we omit them for brevity.